OpenStudy (anonymous):

define f:R->R by f(x)=x^2 +3x -5. use the continuous function definition to prove that f is continuous at 3

OpenStudy (anonymous):

Are we supposed to use the \(\large \epsilon-\delta\) definition or the limit definition of a continuous function?

OpenStudy (anonymous):

Let f:D->R and let c be an element of D then f is continuous at c if for every epsilon>0 there exists a delta>0 such that If(x)-f(c)I<epsilon whenever Ix-cI<delta and x is in D

OpenStudy (anonymous):

Use the fact that \[ f(x)-f(3)=x^2+3 x-18=(x-3) (x+6) \]

OpenStudy (anonymous):

Let us suppose that x< 4, so \[ |f(x)-f(3)|=|(x-3) (x+6)| < 10|x-3| \]

OpenStudy (anonymous):

can I make x less than any number? or does it have to be 4?

OpenStudy (anonymous):

Take \( \delta < \frac {\epsilon }{10} \) If \[ |x-3| < \delta \implies |f(x)-f(3)| < 10|x-3| < 10 \frac {\epsilon}{10}=\epsilon \]

OpenStudy (anonymous):

Your x is going to be near 3 so I took a number a bit bigger than it.

OpenStudy (anonymous):

oh so i could also take 2 and make delta<epsilon/8?

OpenStudy (anonymous):

To be precise one should take |x| < 4 and \( | x-3| <\frac{ \epsilon}{10}\)

OpenStudy (anonymous):

I never know what value to suppose while writing proofs

OpenStudy (anonymous):

But thank you so much..